### v2

# American Sign Language Translation ASLP American Sign Language Production

ASLT
American Sign Language Translation
ASLP
American Sign Language Production
BSL
British Sign Language
BLEU
Bilingual Evaluation Understudy
BLSTM
Bidirectional Long Short-Term Memory
BSLCPT
BSL Corpus T
BSLP
Bilingual Sign Language Production
CNN
Convolutional Neural Network
CRF
Conditional Random Field
CSLR
Continuous Sign Language Recognition
CTC
Connectionist Temporal Classification
C4A
Content4All
DL
Deep Learning
DGS
German Sign Language - Deutsche Gebärdensprache
DSGS
Swiss German Sign Language - Deutschschweizer Gebärdensprache
DTW
Dynamic Time Warping
FC
Fully Connected
FF
Feed Forward
GAN
Generative Adversarial Network
GPU
Graphics Processing Unit
GRU
Gated Recurrent Unit
GSLP
German Sign Language Production
G2PT
Gloss to Pose Transformer
HMM
Hidden Markov Model
HPE
Hand Pose Enhancer
ISL
Irish Sign Language
LSTM
Long Short-Term Memory
MHA
Multi-Headed Attention
MTC
Monocular Total Capture
MSE
Mean Squared Error
MSLP
Multilingual Sign Language Production
MDN
Mixture Density Network
mDGS
meineDGS
mDGS-V
meineDGS-Variants
mDGST
meineDGST
mDGST-H
meineDGST-HARD
mDGST-E
meineDGST-EASY
NMT
Neural Machine Translation
NLP
Natural Language Processing
PHOENIX12
RWTH-PHOENIX-Weather-2012
PHOENIX14
RWTH-PHOENIX-Weather-2014
PHOENIX14T
RWTH-PHOENIX-Weather-2014T
POF
Part Orientation Field
PAF
Part Affinity Field
P2TT
Pose to Text Transformer
RELU
Rectified Linear Units
RNN
Recurrent Neural Network
ROUGE
Recall-Oriented Understudy for Gisting Evaluation
SGD
Stochastic Gradient Descent
SLA
Sign Language Assessment
SLR
Sign Language Recognition
SLT
Sign Language Translation
SLP
Sign Language Production
SMT
Statistical Machine Translation
T2GT
Text to Gloss Transformer
T2PT
Text to Pose Transformer
T2P
Text to Pose
T2G2P
Text to Gloss to Pose
TTS
Text to Speech
WER
Word Error Rate
ASLT
American Sign Language Translation
ASLP
American Sign Language Production
BSL
British Sign Language
BLEU
Bilingual Evaluation Understudy
BLSTM
Bidirectional Long Short-Term Memory
BSLCPT
BSL Corpus T
BSLP
Bilingual Sign Language Production
CNN
Convolutional Neural Network
CRF
Conditional Random Field
CSLR
Continuous Sign Language Recognition
CTC
Connectionist Temporal Classification
C4A
Content4All
DL
Deep Learning
DGS
German Sign Language - Deutsche Gebärdensprache
DSGS
Swiss German Sign Language - Deutschschweizer Gebärdensprache
DTW
Dynamic Time Warping
FC
Fully Connected
FF
Feed Forward
GAN
Generative Adversarial Network
GPU
Graphics Processing Unit
GRU
Gated Recurrent Unit
GSLP
German Sign Language Production
G2PT
Gloss to Pose Transformer
HMM
Hidden Markov Model
HPE
Hand Pose Enhancer
ISL
Irish Sign Language
LSTM
Long Short-Term Memory
MHA
Multi-Headed Attention
MTC
Monocular Total Capture
MSE
Mean Squared Error
MSLP
Multilingual Sign Language Production
MDN
Mixture Density Network
mDGS
meineDGS
mDGS-V
meineDGS-Variants
mDGST
meineDGST
mDGST-H
meineDGST-HARD
mDGST-E
meineDGST-EASY
NMT
Neural Machine Translation
NLP
Natural Language Processing
PHOENIX12
RWTH-PHOENIX-Weather-2012
PHOENIX14
RWTH-PHOENIX-Weather-2014
PHOENIX14T
RWTH-PHOENIX-Weather-2014T
POF
Part Orientation Field
PAF
Part Affinity Field
P2TT
Pose to Text Transformer
RELU
Rectified Linear Units
RNN
Recurrent Neural Network
ROUGE
Recall-Oriented Understudy for Gisting Evaluation
SGD
Stochastic Gradient Descent
SLA
Sign Language Assessment
SLR
Sign Language Recognition
SLT
Sign Language Translation
SLP
Sign Language Production
SMT
Statistical Machine Translation
T2GT
Text to Gloss Transformer
T2PT
Text to Pose Transformer
T2P
Text to Pose
T2G2P
Text to Gloss to Pose
TTS
Text to Speech
WER
Word Error Rate
SignLLM: Sign Language Production Large Language Models
Sen Fang1, Lei Wang2,3, Ce Zheng4, Chunyu Sui5, Mingyu Zhao1, Yapeng Tian6, Chen Chen7
1 Rutgers University, 2 Australian National University, 3 Data61/CSIRO, 4 Carnegie Mellon University
5 Columbia University, 6 The University of Texas at Dallas, 7 University of Central Florida
https://signllm.github.io
Abstract
In this paper, we propose SignLLM, a multilingual Sign Language Production (SLP) large language model, which includes two novel multilingual SLP modes MLSF and Prompt2LangGloss that allow sign language gestures generation from query texts input and question-style prompts input respectively. Both modes can use a new RL loss based on reinforcement learning and a new RL module named Priority Learning Channel. These RL components can accelerate the training by enhancing the model’s capability to sample high-quality data. For SignLLM’s training, we introduce Prompt2Sign, a comprehensive multilingual sign language dataset, which builds from public data, including American Sign Language (ASL) and seven others. This dataset standardizes information by extracting pose information from sign language videos into a unified compressed format. We extensively evaluate SignLLM, demonstrating that our model achieves state-of-the-art performance on SLP tasks across eight sign languages.
Figure 1: Overview: (Left) Major components (e.g., Refer to caption, Refer to caption, Refer to caption, etc.) of our Prompt2Sign dataset. Compressed Pose is reprocessed pose data that is suitable for training, we use public sign language videos to produce compressed pose data in our predefined format; (Right) Our proposed SignLLM aims to generate sign language poses for various application scenarios [12, 6, 110, 105].
1 Introduction
Sign Language Production (SLP) aims to synthesize human-like sign avatars from text inputs. Deep learning-based SLP approaches [71, 72, 37, 73, 74] typically involve sequential steps from text to gloss (i.e., a type of textual vocabulary representing gestures or postures), gloss to pose [10, 52], and finally rendering pose videos into more engaging human-like avatar videos. These processes are complex and challenging to simplify, making sign language data acquisition and processing difficult. This challenge has significantly dampened researchers’ enthusiasm and progress over a considerable period, with the majority of studies in the past decade relying on a German sign language (GSL) dataset named PHOENIX14T [24, 45] for Sign Language Production, Recognition, and Translation tasks (SLP, SLR and SLT). Recent work [86, 22, 4] based on the American sign language (ASL) [20, 78] and other lesser-known languages [19, 25, 58, 68, 31, 80] are relatively rare.
The existing mainstream datasets [14, 20] have significantly advanced the field. However, as time progresses, their limitations become increasingly apparent: (1) These existing datasets consist of different format files, including images, Glosses, subtitles, etc. These images are not easy to be directly trained. Due to redundant information in images that makes it difficult for models to learn essential pose information, training video-level SLP becomes particularly challenging. A way to reduce redundant information is to distill the gesture/posture information into text/npy/json for training. But different datasets pose extraction methods [83, 30, 51, 63, 10, 52] are different. This limitation hampers a specific format model’s ability to use data from other sign languages. (2) Manual annotation for gloss is labor-intensive and time-consuming. (3) Because some videos are obtained from professionals and reprocessed into different formats, scaling up the dataset becomes exceedingly challenging. These limitations collectively impede the development and training of more advanced models.
To solve these issues, we introduce Prompt2Sign, a new multilingual dataset focusing on upper body movements in large-scale signers. The dataset overview is shown in Fig. 1 (Left), showcasing prompts, video subtitles, and files containing digital keypoints information. To create this dataset, we first process the videos using OpenPose [8] to standardize pose information in each frame. Storing keypoints information in our predefined compressed format ([Uncaptioned image], as shown in Fig. 1) can reduce redundancy and facilitate training with seq2seq and text2text models. Subsequently, we reduce reliance on manual annotations by auto-creating prompt words to improve cost-effectiveness. Finally, we improve the processing level of automation for the tools, making the tools highly efficient and lightweight, requiring no additional model loading to process data (i.e., solving the difficulty in manual preprocessing and data collection above). Our new Prompt2Sign dataset is sourced from publicly available sign language datasets and videos on the Internet, covering eight sign languages, making it a comprehensive multilingual sign language dataset. More information is in Table 1 and supplementary materials.
Meanwhile, we recognize that existing models [94, 95, 11, 71, 82, 73] need improvement because training models with our new dataset brings new challenges: (1) Different sign language data cannot usually be trained simultaneously due to text-posture correspondence differences in different sign languages. (2) Handling more languages and a larger dataset results in slow and challenging training processes, with downloading, storing, and data loading difficulties. It is necessary to explore high-speed training methods. (3) The existing model structure cannot grasp more languages and understand more complex natural human conversational inputs. So, we need to explore the aspects overlooked by previous studies, such as multilingual SLP, efficient training, and the ability to understand prompts.
To overcome these challenges, we introduce SignLLM, a large multilingual Sign Language Production (SLP) model developed based on our Prompt2Sign dataset. It can produce the sign language representation of eight languages from texts or prompts. Our SignLLM has two distinct modes: (i) Multi-Language Switching Framework (MLSF), which allows multiple sign languages production in parallel by dynamically adding encoder-decoder groups. (ii) Prompt2LangGloss, allowing SignLLM to support static single-set encoder-decoder generation. Fig. 1 (Right) shows our model inputs and outputs, “thank you” is the input of mode (i), and “how to sign ‘thank you’ in ASL?” is the input of mode (ii). Two multilingual SLP modes deal with different use cases: The MLSF is an efficient mode without semantic confusion, like a dictionary/drawer; The Prompt2LangGloss is a user-friendly mode, like a LLM, it aims to understand complex natural language input. They are like two evolutionary branches diverged from the same species. To address the problem of extended training time caused by more languages and a larger dataset, we utilize the Reinforcement Learning (RL) Loss concepts to quantify the quality of each training batch and prioritize valuable batches through the Priority Learning Channel module.
We conduct extensive experiments and detailed ablation studies. The results validate the superior performance of our SignLLM over baseline approaches [94, 95, 11, 71, 82, 73, 21, 42] on the subsets in eight sign languages. The contributions of this paper can be summarized as follows.
  • A comprehensive multilingual sign language dataset, named Prompt2Sign, featuring an expanded vocabulary and covering eight languages, is introduced. This dataset is designed for compatibility with a wide range of seq2seq models.
  • A large multilingual Sign Language Production (SLP) model with two distinct modes—MLSF for handling text query inputs and Prompt2LangGloss for processing question-style prompts—is introduced. Our method, SignLLM, achieves state-of-the-art performance across eight sign languages in SLP tasks.
  • We introduce a novel reinforcement learning-based loss function, along with a functional module named PLC, as a training strategy for sign language models, designed to reduce training time and computational costs.
2 Related Work
Sign Language Production. In recent years, the field of sign language research has primarily focused on Sign Language Recognition (SLR) [17, 28, 44, 40, 16, 46] and Sign Language Translation (SLT) [7, 15, 43, 4] based on deep learning. Due to a lack of suitable, high-quality datasets for ASL, deep learning researchers conducted their research [71, 72, 37, 73, 74] based on a GSL weather theme dataset, released in 2012 [24, 45]. As previously mentioned, the data processing involved in sign language research is highly complex. Even with the release of a large-scale ASL dataset in 2021 [20], work focused on ASL-related themes based on it has not emerged quickly, as existing work is not easily transferable. The situation is worse for minority languages.
Large Language Models. LLMs refer to giant transformer models trained on extensive textual data, exhibit capabilities in understanding natural language and addressing complex tasks [77, 5, 13, 87, 89]. Sign language is a visual language, theoretically different from language models. However, most current work uses text2text and seq2seq models [66, 99, 48, 61], converting key points/dense maps/grid poses into sequences for training, as opposed to directly training images. Hence, viewing the core process of SLP, text2pose, as a language model is justifiable. Extensive research indicates that an increase in parameters or data volume [34, 41] significantly enhances the abilities of LLMs [65, 5, 13]. There are more than a hundred sign languages in the world, most of which have datasets in video form. Therefore, conducting advanced research to address the anticipated surge in data volume in the future is of paramount importance. In this work, we aim to enable the model to generate sign languages across diverse linguistic backgrounds, ensuring its adaptability to our new dataset.
3 Our Benchmark: Prompt2Sign
Existing Datasets Weaknesses. The main shortcoming of previous work is the lack of unified data storage formats, when there is a mismatch between these models in Sign Language Production (SLP) [94, 95, 11, 71, 82, 73] and Sign Language Translation (SLT) [7, 15, 43, 4], it can lead to complex challenges: (1) The results of the SLT model are difficult to use as training data for the SLP model directly due to format incompatibility (e.g., [86, 6] & [70, 75]) (2) The results of the SLP model are difficult to use as input for the SLT model (evaluation experiment needs, e.g., [6] & [4]). (3) The output of the SLP model is not suitable as input for most style transfer models (e.g., [11, 97, 107], the researchers have to train a pose2video model themselves). Therefore, we create a standardized dataset to address data collection, utilization, and storage challenges.
Data Collection. Our data collection process, in contrast to previous methods, includes the following steps: (1) downloading sign language videos in specific languages from the Internet and public datasets [19, 25, 58, 68, 31, 80]; (2) editing and aligning these videos; (3) extracting 2D keypoints from each video frame using OpenPose [8] and saving them as JSON files; (4) calculating the 3D poses, which are stored in a predefined compressed data format. See further details below and in the supplementary materials.
Dataset Range. We have cited the sources of our publicly available data, and some of the more popular works were not considered due to their limited accessibility and potential usage restrictions. Additionally, while there are other multilingual datasets available [29, 101, 55, 33], they may not possess the same level of comprehensiveness as ours. For example, [29] and [101] translate two types of sign language videos into spoken language (SLT), while our work is from spoken language to videos (SLP). We aim to establish a robust multilingual SLP method with a dataset that supports a broader range of application scenarios. For example, [29] is limited to Bible translation, and [101] focuses solely on cross-lingual SLT. In contrast, our work covers a broader range of scenarios and comprehensively addresses the SLP task. Additionally, some datasets, such as [55] and [33], are restricted to multilingual dictionary forms.
Subset ASL GSL DSGS LSF-CH LIS-CH LSA KSL TSL
Train 31,047 7,096 8,043 5,672 2,254 2,400 700 28,142
Dev 1,739 519 500 500 250 400 300 4,418
Test 2,343 642 500 250 250 400 200 3,742
Table 1: Dataset Statistics [19, 25, 58, 68, 31, 80, 57, 100] : The number of video clips in their train, dev, test set, respectively. Languages included: American (ASL), German (GSL, Alias DGS), Swiss German (DSGS), French Sign Language of Switzerland (LSF-CH), Italian Sign Language of Switzerland (LIS-CH), Argentine (Lengua de Señas Argentina, LSA), Korean (KSL), and Turkish (TSL). “Sign Language” is omitted in some full names.
Figure 2: (Left) MLSF contains parallel Enc-Dec groups (i.e., Text2Pose $\times$ number of languages), the Prompt2LangGloss adds a language attribute marker at the gloss channel (i.e., Text2Gloss2Pose $\rightarrow$ Prompt2LangGloss2Pose). (Right) The output of SignLLM can be converted into a skeletal pose video, which can then be rendered into a realistic human appearance by vid2vid models [69, 11, 97, 107, 105].
Unified Compression of Pose Data. Building on previous work [104, 71, 75], we develop a three-step tool for standardizing data processing. The tool is highly efficient and lightweight, requiring no additional model loading and supporting large-scale data processing. Furthermore, we optimize it specifically for sign language data processing (e.g., removing unnecessary leg movement computations and integrating these optimizations into Prompt2Sign’s pipeline). The main steps of Data Collection can be visualized as: [Uncaptioned image] to [Uncaptioned image] to [Uncaptioned image]. Among all the steps, the most crucial part is the transition from 2D to 3D compressed pose data:
  • Step I: First, we obtain the length of the skeleton through the 2D keypoint coordinates ($x$ and $y$), $a$ and $b$ represent indices that identify the two keypoints (or joints) forming a bone $L=\sqrt{(ax-bx)^{2}+(ay-by)^{2}}$.
  • Step II: We compute the 3D rotation angles from 2D keypoints data: $A_{x},A_{y},A_{z}=\frac{\text{angle}_{x},\text{angle}_{y},\text{angle}_{z}}{\sqrt{\text{angle}_{x}^{2}+\text{angle}_{y}^{2}+\text{angle}_{z}^{2}}}$, where $A$ represents the normalized angles.
  • Step III: We define $P_{x},P_{y},P_{z}$ as the starting joint coordinates computed from 2D keypoints, and define $Q_{x},Q_{y},Q_{z}$ as the target coordinates in 3D space: $Q_{x}=P_{x}+L\times A_{x}$, $Q_{y}=P_{y}+L\times A_{y}$, $Q_{z}=P_{z}+L\times A_{z}$.
These mathematical formulas initialize the skeletal model by calculating skeletal length $L$, root node position, rotation angle, and 3D coordinates ($x$, $y$, $z$), serving as input for simulating 3D human skeletal motion. Compared to previous methods [104], this approach is much more efficient, focusing solely on extracting the posture and gesture information relevant to sign language from the video data. This preprocessing step discards redundant information, reduces data size by 80% relative to the raw video, and standardizes the format, enabling easier integration with text-to-text and sequence-to-sequence models and applications without the need for sign language-specific data loaders.
Dataset Statistics. After the data processing, train, dev, and test sets of different language parts are shown in Table 1. We constructed 120 English templates and 210 prompt word templates for other languages (with 30 templates for each language), which are randomly associated with the oral text data to form the prompt part. The templates are carefully selected from hundreds of sentences generated by LLM that cover most everyday expressions.
4 Our Model: SignLLM
4.1 Preliminary of Text2Pose Method
The general SLP pipeline (i.e., text to sign language video) [71, 72, 37, 73, 74] has following steps: text-to-gloss conversion, gloss-to-pose mapping, and finally pose-to-video rendering. In our work, we mainly focus on the first two steps.
Text2Gloss & Gloss2Pose. Essentially, the transformation from text-to-gloss and gloss-to-pose can be distilled to a sequence-to-sequence [71, 72] problem in the realm of textual data, and their structures to bear significant resemblances. We define $x_{u}$ as the input text $x$ tokens at position $u$ (total number is $\mathcal{U}$, position from 1 to $\mathcal{U}$), $p_{w}$ as the output pose $p$ at position $w$ (total number is $\mathcal{W}$, frame position from 1 to $W$), and then we use an encoder-decoder transformer framework to convert input into output:
$f_{u}=Enc_{input2output}(x_{u}|x_{1:\mathcal{U}})$ (1)
$p_{w+1}=Dec_{input2output}(p_{w}|p_{1:w-1},f_{1:\mathcal{U}})$ (2)
Here, $f_{u}$ denotes the encoded source of $x_{u}$. The output text tokens generated from this process form the input for the next stage of our translation model. In short, text2pose is the core method of SLP, and some researchers use gloss as an intermediate to make it text2gloss2pose.
4.2 Design Overview
Architecture. SignLLM has two modes Multi-Language Switching Framework (MLSF) and Prompt2LangGloss, as shown in Fig. 2 (Left), both make the model capable of multilingual sign language production by using the multilingual Prompt2Sign dataset. Both modes can be trained by new RL Loss and Priority Learning Channel module.
Motivation. Unlike existing SLP models, LLM directly fine-tuning for text-to-pose translation would hamper its dialogue capabilities, while using it without fine-tuning would treat translation requests as questions rather than performing actual text-to-pose translations. Therefore, we propose two specialized modes: MLSF for direct text-to-pose translation and Prompt2LangGloss for LLM-based interaction (e.g., “how to sign ‘thank you’ in ASL?” as input). These modes, analogous to evolutionary branches of the same species, serve complementary purposes in multilingual SLP.
4.3 Two Multilingual SLP Modes
MLSF is a mode that most existing models can refer to, like a dictionary/translator, and Prompt2LangGloss mode is designed specifically for sign language LLM. They are like two evolutionary branches diverged from the same species.
Multi-Language Switching Framework. It can be understood as having multiple parallel Text2Pose channels/Enc-Dec groups, each language has an Enc-Dec group, allowing each channel/Enc-Dec to be independently trained and inferred. Text2Pose visual representation is shown on the left of Fig. 2, the red rectangle represents the eight Enc-Dec in our model, and the middle partition represents the parameters stored separately in different Enc-Dec groups. The assignment operation could be formalized as $\text{Enc}_{\mathcal{L}}=\mathcal{E}_{\mathcal{L}}$ and $\text{Dec}_{\mathcal{L}}=\mathcal{D}_{\mathcal{L}}$. Here, $\mathcal{L}$ denotes the language of input, while $\mathcal{E}_{\mathcal{L}}$ and $\mathcal{D}_{\mathcal{L}}$ are the mapping from language $\mathcal{L}$ to an encoder and decoder in the sets $\mathcal{E}$ and $\mathcal{D}$. Similar to selecting tools from a drawer, MLSF allows you to choose the appropriate $\mathcal{E}_{ASL}$-$\mathcal{D}_{ASL}$ pair from the $\mathcal{E}$-$\mathcal{D}$ groups for training and inference. This modular design functions like a language drawer system, where each language Enc-Dec component can be accessed and utilized on demand from eight pairs.
Prompt2LangGloss. While MLSF tackles multilingual support through architectural design, Prompt2LangGloss approaches the challenge from a linguistic perspective. This mode introduces a novel intermediate representation that bridges the gap between natural language understanding and sign language generation. Enriching the traditional gloss notation with language-specific attributes creates a more nuanced and contextually aware translation process.
Gloss, essentially a shorter textual representation of sign language gestures, operates as an intermediate entity when using a text2pose model. As shown in Fig. 2 (Left), our proposed enhancement of this model involves appending an additional language attribute to each text word during the reading and tokenizing stages. For instance, a traditional gloss token “<xxx>" can be transformed into “<ASL_xxx>", thus introducing a LangGloss layer of conditional input $f_{u}=Enc_{t2lg}(x_{u}|x_{1:\mathcal{U}})$ into SLP based on Eq. (1): $lg_{w+1}=Dec_{t2lg}(lg_{w}|lg_{1:w-1},f_{1:\mathcal{U}})$. So, our LangGloss is a pseudo-Gloss used to distinguish language information in the parameter space by identifying language attributes. This way, we solve several challenges: (1) LangGloss allows the existing models to train multilingual data. By adding language attributes to gloss, we reduce the semantic ambiguity that occurs when words share the same form across different languages. (2) LangGloss, as a mediator, can solve the limitations of the existing model in understanding complex, natural human inputs. It reduces the negative impact of directly processing intricate prompt words, improving the model’s response accuracy.
In short, these two modes create a robust multilingual SLP system: This dual-mode design enables SignLLM to achieve processing efficiency and semantic accuracy in existing models and sign language LLMs, respectively.
4.4 Reinforcement Learning Training Strategy
To reduce training time, we utilize the RL reward concept to quantify each training batch’s quality and prioritize valuable batches through the Priority Learning Channel module. Both Multilingual SLP modes can use RL Loss and PLC.
Reinforcement Learning Loss. Reinforcement Learning (RL) has an important advantage in identifying high-value actions or samples, allowing for prioritized learning of valuable data. This approach could help address the challenge of slow training when using more sign languages. But before learning high-value data batch, we should transform the ordinary generative model first into an RL-like model, so we design RL Loss for model transformation.
Figure 3: RL elements: User, Agent, Environment, Cyclic Sampling, PLC to sketch the sequence prediction learning process.
Concretely, we set the input sequence as the state $s_{t}$, and the output sequence is the action $a_{t}$, and the reward $r_{t}$. $t$ stands for time, $i$ stands sample, $\hat{y}_{i}$ represents the model’s predicted output, and $N$ denotes the total number of samples. The closer the prediction is to reality (mean squared error), the greater the reward: $r=-\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$. With this interpretation, we can reformulate the traditional supervised learning problem of minimizing MSE loss to maximize the expected cumulative reward, where $L$ denotes the MSE loss function, $M$ is the model, and $x_{t}$, $y_{t}$ are the model inputs and corresponding targets respectively:
[…]
where $\theta$ represents the trainable parameters of the model, $E_{\theta}$ denotes the expected value with respect to the model parameters $\theta$, and $T$ is the total number of time steps in the sequence. Here, $\operatorname{argmax}$/$\operatorname{argmin}$ returns the value of $\theta$ that maximizes/minimizes the expected cumulative reward. These optimized parameters $\theta$ are founded by using gradient descent, updating parameters proportionally to the gradient of expected cumulative reward concerning model parameters. In summary, we quantify the effectiveness of MSE optimization at each time step as reward $r$. Then the $r$ will serve as the input for our Priority Learning Channel.
Priority Learning Channel. The RL Loss itself does not possess subjective acceleration capabilities, it is designed for PLC to prioritize the learning of more valuable data. We have defined rewards $r$, sample $i$, and data for each batch $j$. They then are converted into sampling probabilities for each data sample according to $P(i)=\frac{r(i)^{\eta}}{\sum_{j\in S}r(j)^{\eta}}$, where $\eta$ regulates the intensity of prioritization, and $S$ represents the dataset. By employing these sampling probabilities, the choice of data samples for each batch is no longer uniform but regulated by their respective rewards (e.g., if the reward is less than 50%, skip the batch). The per-step RL loss, $L(i)$, is computed for the chosen data, which is then used to optimize the model parameters following the policy gradient theorem. This procedure is formally expressed as $Minimize\;\;E_{i\sim P(i)}[L(i)]$ (where $E$ is the expectation), the whole RL Loss and PLC system is shown in Fig. 3. By continually updating the model based on the most rewarding samples, the PLC brings the advantages of RL to sequence prediction tasks. The adaptive nature of the PLC ensures that the model’s focus shifts by the model’s evolving knowledge, thereby accelerating the learning process.
DEV SET TEST SET
Type: BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE
NSLP-G [38] - - - - - 5.75 8.21 11.62 17.55 31.98
Fast-SLP Transformers [22] 17.19 23.11 29.49 36.96 55.85 12.85 17.35 23.38 39.46 46.89
Neural Sign Actors [2] - - - - - 13.12 18.25 25.44 41.31 47.55
SignLLM-1x40M-Base-M (ASL) 18.77 25.42 32.44 40.66 61.44 14.13 19.08 25.72 43.40 51.57
SignLLM-1x120M-Large-M (ASL) 19.40 26.11 33.38 41.89 63.21 14.52 19.65 26.44 44.72 53.14
SignLLM-1x1B-Super-M (ASL)  20.09 +2.90 27.04 +3.93 34.45 +4.96 43.16 +6.20 65.19 +9.36  15.03 +2.18 20.28 +2.93 27.35 +3.97 46.17 +6.71 54.86 +7.97
SignLLM-1x40M-Base-P (ASL) 17.34 23.57 29.87 37.81 56.93 13.06 17.66 23.77 40.15 47.76
SignLLM-1x120M-Large-P (ASL) 18.05 24.28 31.04 38.97 58.78 13.48 18.27 24.57 41.57 49.42
SignLLM-1x1B-Super-P (ASL)  18.68 +1.49 25.11 +2.00 31.99 +2.50 40.14 +7.18 60.47 +4.62  13.93 +0.92 18.86 +1.51 25.40 +2.02 42.87 +3.41 50.91 +4.02
Table 2: American Sign Language Production (ASLP): Comparison of SignLLM variants with baseline on Text to Pose task by using our Prompt2Sign ASL part. The improvement (+num) is relative to the latest work [22].
DEV SET TEST SET
Type: Language: BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE
SignLLM-6x40M-Base-M DSGS 9.73 15.82 19.85 24.84 37.57 7.34 9.89 16.85 26.04 31.51
SignLLM-6x120M-Large-M DSGS 11.45 17.28 19.84 29.07 41.69 9.69 14.06 16.35 29.80 31.51
SignLLM-6x40M-Base-M LSF-CH 9.79 18.48 23.13 28.86 34.98 8.92 13.12 16.11 26.48 37.92
SignLLM-6x120M-Large-M LSF-CH 13.72 20.79 23.40 25.15 38.39 9.60 12.58 16.98 22.71 41.96
SignLLM-6x40M-Base-M LIS-CH 10.81 14.46 19.93 24.55 35.83 7.34 10.56 15.24 22.73 36.42
SignLLM-6x120M-Large-M LIS-CH 12.10 18.04 23.01 25.95 36.98 9.30 11.20 15.68 23.38 38.37
SignLLM-6x40M-Base-M LSA 10.72 15.55 21.76 25.91 38.78 7.33 14.86 16.68 22.55 34.42
SignLLM-6x120M-Large-M LSA 11.69 14.79 26.25 28.08 39.01 8.21 11.04 17.05 26.68 37.46
SignLLM-6x40M-Base-M KSL 9.42 14.67 17.24 26.41 31.96 8.31 11.84 17.93 24.15 33.78
SignLLM-6x120M-Large-M KSL 12.91 19.45 15.17 24.07 37.83 10.09 13.06 18.37 25.75 33.69
Hybrid Translation System [42] TSL - - - - - 12.64 18.28 31.48 53.17 -
SignLLM-6x40M-Base-M TSL 14.53 19.86 29.93 36.86 58.01 13.23 17.80 25.39 39.30 57.03
SignLLM-6x120M-Large-M TSL 15.17 21.70 31.73 38.86 71.10 14.36 18.74 26.96 43.21 57.12
Table 3: Multilingual Sign Language Production (MSLP): Comparison of different SignLLM M-mode variants with a baseline on Text to Pose task. We propose the first Multilingual SLP benchmark, with the exception of the existing TSL-Baseline.
5 Experiments and Discussions
Setup. We provide the naming rules as follows: SignLLM-{number of languages}x{single language parameters}-{submode size}-{the mode of training}-{the language of input}. Such as “SignLLM-2x40M-Base-M (ASL)”, the nomenclature “2x40” denotes that the model comprises 2 language knowledge, with each language component estimated to be around 40 million parameters in size, and a total is 80 million parameters (“1B” represents a total of 1 billion parameters). There are Base, Large, and Super versions, depending on a single language parameter size provided by the model. The encoder and decoder of our model versions (i.e., Base, Large, Super) both have two layers. When the model is expanded to Large and Super versions, the layers are unchanged, the parameters are expanded by about two and four times, respectively. M and P stand for models trained using MLSF and Prompt2LangGloss. At the end is the language of the input model, ASL, GSL, LSA etc.
DEV SET TEST SET
Approach: BLEU-4 ROUGE BLEU-4 ROUGE
Progressive Transformers [71] 11.82 33.18 10.51 32.46
Adversarial Training [70] 12.65 33.68 10.81 32.74
Mixture Density Networks [72] 11.54 33.40 11.68 33.19
Mixture of Motion Primitives [73] 14.03 37.76 13.30 36.77
Photo-realistic SLP [75] 16.92 35.74 21.10 42.57
Fast-SLP Transformers [22] 18.26 39.62 22.15 46.82
SignLLM-1x40M-Base-M (GSL) 18.61 40.69 22.76 48.05
SignLLM-1x120M-Large-M (GSL)  19.31 +1.05 41.42  23.25 +1.10 49.08
SignLLM-1x1B-Super-M (GSL) 19.07 41.83 +2.21 23.21 49.52 +2.70
SignLLM-1x40M-Base-P (GSL) 17.12 37.43 20.93 44.21
SignLLM-1x120M-Large-P (GSL) 17.55 38.10 21.39 45.16
SignLLM-1x1B-Super-P (GSL) 17.54 38.48 21.35 45.57
Table 4: German Sign Language Production (SLP): Comparison of different models with existing work on Text to Pose task. The improvement (+num) is relative to the latest work [22].
Metrics. We mainly use 3 metrics: (i) BLEU-n score measures the similarity between machine-generated translations and reference translations based on n-grams, the closer the predicted result is to the input (reference), the higher the value. BLEU-n [60] means that n words are used as the basic computing unit, and the higher the n, the higher the fluency requirement. (ii) ROUGE score [49] is similar to BLEU, but it is more concerned with consistency and coverage. It indicates better agreement between the generated and reference texts, indicating a more accurate and comprehensive summary. (iii) DTW score underpinned by dynamic programming principles [3], is employed to ascertain the smallest manipulation distance between clips and sentences; the lower, the better. More model information and experiments can be found in supplementary materials.
DEV SET TEST SET
Approach: BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE
Base 10.96 14.68 22.49 40.29 44.13 9.27 10.72 18.64 39.88 40.39
Prompt2LangGloss 15.78 22.66 30.27 48.30 50.88 16.32 20.27 28.70 47.89 48.18
RL Loss 18.33 24.28 32.18 48.83 52.61 13.26 17.33 24.51 40.76 40.95
RL Loss & PLC 18.77 25.42 32.44 52.66 61.44 14.13 19.08 25.72 43.40 51.57
Table 5: Comparison of Different Modules: SignLLM-40M-Base results for Text to Pose task on the ASL part of Prompt2Sign, with multiple data augmentation techniques. Base: MLSF with Normal MSE Loss. Prompt2LangGloss: Prompt2LangGloss with normal MSE Loss. PLC: Priority Learning Channel. Top performances are highlighted in bold, while second top performances are underlined.
5.1 Quantitative Evaluation
Back Translation. Back translation means translating generated sign language videos back into spoken language sentences. These sentences are compared with the original input sentences to evaluate the translation quality. The task is widely adopted to evaluate Sign Language Production (SLP) as it can indicate the accuracy of the produced sign language videos [71], translation models are trained on the corresponding language skeletal pose data.
In Table 2, we conduct American Sign Language Production back-translation tests using SignLLM on the ASL part of our new dataset, and Table 4 further compares our method with other recent approaches for German SLP on the GSL part of Prompt2Sign dataset [71, 70, 72, 73, 22]. These two languages stand for high-resource languages (i.e., languages with rich data resources), and our comprehensive tests at different levels on the dataset demonstrate impressive performance compared with the latest works in the field. The results affirm the competitiveness and potential of our proposed method, regardless of the specific sign language in use. More evaluation results and analyses for these two languages can be found in the supplementary materials.
In Table 3, we present the results of our model for six different sign languages. These six languages represent low-resource languages, and they are considered limited languages. For languages with low resources, their vocabulary, video time, and diverse corpus sources are relatively low, making training more difficult. From the table data, it can be observed that our performance remains strong in languages where training data is lacking. As long as the input text/prompt can be encoded as a computationally recognizable word and video exists, our method can translate it into the corresponding language pose video after training.
5.2 Ablation Evaluation
Figure 4: RL Training Efficiency Analysis: Comparison of different Settings on DTW values (the lower the better) at different epoch times. Left Y-axis: Value of DTW. Right Y-axis: Value of Loss. Prompt: Prompt2LangGloss mode.
Performance. Table 5 indicates our last two innovative strategies (RL Loss and PLC module) significantly improve the model’s performance, and both strategies contribute to substantial improvements. We find that with sufficient data, LangGloss can outperform the base model without enhancements, showing good usability. So, we conducted more research on this in the supplementary materials. In addition, despite the demonstrable improvement, RL Loss facilitates in the development set (showing improvements of 7.37 in BLEU-4 and 8.48 in ROUGE scores), its effectiveness is less apparent in the test set. However, a completely different picture emerges when we combine RL Loss with the PLC. This innovative integration improves the ROUGE score by 25.9% on the test set (from 40.95 to 51.57), showcasing the positive impact of our proposed technique. These results show that PLC has successfully enhanced the adaptability of the original RL Loss to unknown environments.
In Fig. 4, we compare the Base model’s performance with different module schemes. The ablation comparison is primarily based on observing the variability of DTW scores across epochs to assess the effectiveness of each approach. We observe that: (i) The standalone use of the Prompt2LangGloss model exhibits the lowest efficiency, as it introduces noise by incorporating prompts and tokenizers, which is also evident in Table 5. (ii) The combination of the two RL methods shows modest performance improvement compared to using each method individually, although it is not statistically significant. Future work could explore adaptive weighting mechanisms or hierarchical integration of these methods to leverage their complementary strengths fully. (iii) In terms of training efficiency, they significantly reduce training time. Compared to the base setting, our RL Loss with PLC approach reduces training time by 27.1%, which is crucial for training large-scale sign language models. Our ablation studies reveal that RL-based sample prioritization not only enhances training efficiency but also provides a scalable, unified, multilingual SLP framework.
Figure 5: We use an adjusted vid2vid model [23] to convert the predicted skeletal pose video into a more realistic final video.
5.3 Qualitative Evaluation
Qualitative Presentation. We use our predicted pose results as input, which are then used to generate rendered videos in Fig. 5. We can observe that our video outcomes are of high quality, with highly accurate finger movements and high image fidelity. Our results surpass all previous works, benefiting not only from technological advancements but also from the superior output quality of our SignLLM compared to previous smaller models: the postures we predict are rarely missing, unlike previous works that often suffered from issues such as flickering, incomplete or missing fingers, and low input quality due to densely packed fingers, so it works well with the latest model [23]. However, the final sign language video finger problem still exists anyway, and in the future, we can consider the special optimization of these style transfer vid2vid models.
DEV SET TEST SET
Approach: BLEU-4 ROUGE BLEU-4 ROUGE
Progressive Transformers [71] 10.79 36.15 9.59 35.42
Fast-SLP Transformers [22] 16.68 43.2 24.24 51.05
SignLLM-1x40M-Base-M (GSL) 16.96 44.41 24.74 52.45
SignLLM-1x120M-Large-M (GSL)  17.73 +1.05 45.11  25.39 +1.15 53.45
SignLLM-1x40M-Base-P (GSL) 16.27 43.76 24.12 54.35 +3.30
SignLLM-1x120M-Large-P (GSL) 16.71 45.40 +2.20 25.04 52.80
Table 6: Presentation Effect Study: Results of Text to Sign task in GSL. The improvement (+num) is relative to the latest work [22].
In Table 6, we conducted a series of final video back-translation evaluations (as shown in Fig. 5) based on the German SLP task to investigate two main research questions: (i) Whether rendering the predicted results into real sign language videos would lead to a decrease in accuracy. (ii) How does our work compare to previous studies on the task of text-to-sign-language real video generation? Based on our observations, there was generally not a significant accuracy loss, but there were some fluctuations compared to the baseline. Our approach outperformed previous works, which could be attributed to the higher quality of our data, making it more suitable as input for style transfer models.
5.4 Discussion
Societal Impact. Our model has the potential to assist people with disabilities in three key areas: sign language teaching, generative sign language translation, and real-time interpretation for broadcasting. (1) Traditional sign language teaching relies heavily on human instructors and static pictorial representations, limiting learning accessibility. (2) Current sign language translation software remains inadequate for effective communication between deaf people and their family members who lack sign language knowledge. (3) Additionally, real-time sign language interpretation is only available for limited content like major news broadcasts, creating barriers for the deaf community in daily life. However, its current level of accuracy is not high enough to be fully trusted, and users must be cautious to use it.
Limitation. Our tool enhances sign language data processing automation but isn’t fully end-to-end. Manual preprocessing is still required for OpenPose processing, video editing, and transcript alignment, etc. For pose video to final video conversion, when we use a style transfer model to make the final video, it requires more complex processing of the pose video. So there is still a way to go before large-scale use, which requires us to make industrial-level adaptation improvements to the output style of SignLLM.
6 Conclusion
We present SignLLM, a large multilingual SLP model. For training this model, we propose Prompt2Sign, a standardized dataset that contains eight sign languages. Our model with two modes, MLSF and Prompt2LangGloss, progressively incorporates more sign languages while preserving translation efficiency and introducing LLM capabilities. Our new RL loss and new PLC module solve the challenge of longer training time due to more data. Finally, we show baseline comparisons, ablation studies, experiments under various parameters, and qualitative evaluations for discussion, which proves the efficacy of our methodology.